Copper has a density of 9 g/cm3 and a bulk modulus of 1.2 x 1011 N/m2. What isthe speed of a wave through it?


1
Expert's answer
2022-03-06T15:16:55-0500

The speed of wave is given as follows:


"c = \sqrt{\dfrac{K}{\rho}}"

where "K = 1.2\times 10^{11}N/m^2" is the bulk modulus, and "\rho = 9g/cm^3 = 9000kg/m^3" is the density. Thus, obtain:


"c = \sqrt{\dfrac{1.2\times 10^{11}N/m^2}{9000kg/m^3}} \approx 4\times 10^3m/s"

Answer. "4\times 10^3m/s".